OpenStudy (anonymous):

Solve the equation 4^2x+3=1

OpenStudy (helder_edwin):

is it \[\large 4^{2x}+3=1 \] or \[\large 4^2x+3=1 \] ?

OpenStudy (anonymous):

2x+3 is an exponent

OpenStudy (helder_edwin):

so it is \[\large 4^{2x+3}=1 \]

OpenStudy (anonymous):

yes :)

OpenStudy (helder_edwin):

u can write like this \[\large 4^{2x+3}=1=4^0 \] from which u get \[\large 2x+3=0 \]

OpenStudy (helder_edwin):

and from this x=-3/2

OpenStudy (anonymous):

but how did you turn 1 into 4^0

OpenStudy (helder_edwin):

because 4^0=1.

OpenStudy (anonymous):

ohhh.. okiee thanks sooo muchh!!

OpenStudy (helder_edwin):

u r welcome